The primary energy consumption in Germany increased from 2,000 Mtoe in 2012 to 2,300 Mtoe in 2013. The GDP increased from 15,197 billion Euros in 2012 to 16,603 billion Euros in 2013 at constant 2,100 prices. What was the GDP elasticity of energy demand in Germany?

b) Discuss in details the number of difficulties that could be experienced when analyzing the energy demand of developing countries.
Answer #1

Price elasticity of energy demand=

First calculate the elasticity of energy consumption= (2300-2000)/ 2000=0.15

Then calculate elasticity of GDP=(16603-15197)/15197=0.09

Finally elasticity= 0.15/0.09=1.67

b) It is a known fact that the developed economies consume more energy than its developing counterparts.There is a lot of energy lost at every stage in the developing economies. However it is difficult to analyse the energy demand of developing countries because:

1. There is no sufficient data available about the energy consumed. It is still possible to understand the energy consumption in the production sector but at household level it becomes difficult.

2. There is no data available between the trading activities of fuel or its by products between neighboring countries.